beauregard jingles has $2.05 in dimes and nickles, including five more nickels than dimes. How many dimes does Beau have?
Answer by nerdybill(7384)

beauregard jingles has $2.05 in dimes and nickles, including five more nickels than dimes. How many dimes does Beau have?
.
Let d = number of dimes
then
d+5 = number of nickels
.
.10d + .05(d+5) = 2.05
.10d + .05d+.25 = 2.05
.15d+.25 = 2.05
.15d = 1.80
d = 12 (dimes)
.
nickels:
d+5 = 12+5 = 17